OpenStudy (anonymous):

How many five-number combinations can be made from the following five numbers? 55111

OpenStudy (raden):

use the formula : n!/(a!b!) here, n=5, a=2, b=3

OpenStudy (anonymous):

im only should be 6th but they skipped me to the 9th grade this stuff is confusing to me

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

so you dont know about factorial ? the other way is by manual, (but it too long for other case)

OpenStudy (anonymous):

@RadEn no i just need help i thought it was 30

OpenStudy (raden):

well, you can finish it. it is 5!/(2!3!) = 5.4.3.2.1/(2.1.3.2.1) = 120/12 = ... ?